Title

Article text.

INAF queue observing, 2016 December 6 - 7

Observer: R. Carini, A. Marchetti
Telescope Operator: S. Allanson
Support Astronomer: B. Rothberg on call
Instrument: MODS1, MODS2, LUCI1

Summary

LUCI2 is not available

At sunset there is strong southwest wind and some clouds.

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:45-12.45 UT) = 11h 50m

Observing time: 5h 02m (43%)

Weather time loss: 6h 48m (57%)

technical time loss 0m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Pancino H12 Dual 0.5 h COMPLETED MODS1+MODS2
Massaro 3FGL60258.2+3555 prism, Red 0.6 h (300sec x6) COMPLETED MODS1+MODS2
Pancino H18 Dual 0.5 h (250sec x6 in each filter) COMPLETED MODS1+MODS2
Pancino H3 Dual 36 m (1100sec x2 in each filter) filter MODS1+MODS2,
Calibrations Annibali MODS1 flats and arcs for DD068_RunB Dual    
Calibrations Antoniucci MODS2 flats, arcs, imaging flats Dual    
Calibrations Massaro MODS1-MODS2 bias 3kx3k, 4kx4k Red    
Calibrations Massaro MODS1-MODS2 flats, arcs 1", 0.3" Red    
Standard Pancino MODS1-MODS2 g191b2b Dual   seeing oscillating, some clouds
Annibali MODS1-MODS2 DD068_RunB Dual 1200sec x2 parallactic angle out of spec
Annibali MODS1-MODS2 Feige34 Dual   parallactic angle out of spec

Details

01:01 we send the coordinates of Pancino MODS target H12

01:18 the TO is collimating

01:24 we send the acquisition for Pancino MODS target H12

01:40 seeing is between 1.3" and 1.6"

01:41 we take acquisition images

  • mods1r.20161207.0002-0004

  • mods2r.20161207.0002-0004

01:46 we start the science of Pancino MODS target H12 (binocular)

  • mods1r.20161207.0005-0007

  • mods1b.20161207.0002-0004

  • mods2r.20161207.0005-0008

  • mods2b.20161207.0002-0004

01:55 seeing 1.2"

02:02 seeing 1.6"

02:03 seeing is 1.8"

02:05 seeing has jumped to 2.3". We acquire target of Massaro_3FGL60258.2+3555

02:08 seeing is 2". Acquisition:

  • mods1r.20161207.0008-0010

  • mods2r.20161207.0009-0011

07:28 we send the science of Massaro. Seeing is 1.5", but keeps oscillating

  • mods1r.20161207.0011-0013

  • mods2r.20161207.0012-0014

02:48 seeing 0.9". We slew to Pancino H18. Acquisition:

  • mods1r.20161207.0014-0017

  • mods2r.20161207.0015-0017

03:06 seeing is 0.7". We send the science of Pancino H18:

  • mods1r.20161207.0018-0020

  • mods1b.20161207.0005-0007

  • mods2r.20161207.0018-0020

  • mods2b.20161207.0005-0007

03:27 seeing is 1.5". We slew to Pancino H3.

03:30 the TO says there's no guiding star. He has to check the pointing of the telescope. MODS1 says: preset IIF error. MODS2 is going on with the acquisition.

03:34 we start the script again. Acquisition

  • mods1r.20161207.0021-0025

  • mods2r.20161207.0022-0024 (0021 is BAD)

03:38 strong wind gusts

03:56 we send the science of Pancino H3. Seeing is 1"-1.2".

04:04 the seeing is getting worse, peaks of 1.5"-1.6"

04:08 seeing is variable between 1" and 1.4"

04:09 clouds are becoming thick and wind is raising too much

04:14 guide star lost

04:17 we stop the integration.

  • mods1r.20161207.0026

  • mods1b.20161207.0008

  • mods2r.20161207.0025

  • mods2b.20161207.0008

  • each image is about 1100 sec instead of 1800

04:18 the TO closes the dome

06:42 We start calibrations

for Annibali program DD068_runB.cal (flats and arcs)

  • mods1r.20161207.0027-0031

  • mods1b.20161207.0009-0013

for Antoniucci MODS2 (arcs)

  • mods2r.20161207.0026-0028

  • mods2b.20161207.0009-0011

for Antoniucci MODS2 (flats)

  • mods2r.20161207.0029-0034

  • mods2b.20161207.0012-0017

for Antoniucci MODS2 (red imaging flats)

  • mods2r.20161207.0035-0039

  • mods2b.20161207.0018-0023

for Massaro (bias 3kx3k)

  • mods1r.20161207.0032-0041

  • mods1b.20161207.0014-023

  • mods2r.20161207.0040-0049

  • mods2b.20161207.0023-0032

for Massaro (bias 4kx3k)

  • mods2r.20161207.0050-0060

  • mods2b.20161207.0033-0048

  • mods1r.20161207.0042-0052

  • mods1b.20161207.0024-0039

for Massaro (red flats). These are close to SATURATION. We will repeat them tomorrow

  • mods1r.20161207.0053-0057

  • mods2r.20161207.0061-0065

for Massaro (1" red arcs). The PI choose 1" for the comparison lamps, but Olga says it's far too large, we need to use 0.3"

  • mods1r.20161207.0058-0059

  • mods2r.20161207.0066-0067

for Massaro (0.3" arcs)

  • mods1r.20161207.0060-0063

  • mods2r.20161207.0068-0071

  • mods1b.20161207.0040-0041

  • mods2b.20161207.0049-0050

this script took also lamps for the blue channel

10:00 the TO opens the dome

10:01 we send the coordinates of the standard for Massaro. It goes in the direction of the wind, we change target.

10:17 we send the coordinated for Pancino spect-phot standard, G191B2B

10:25 the sky is getting covered again, but seeing is 0.9".

10:29 the telecope lost the guide star and found it again

10:31 seeing is 1.4"

10:36 we send the acquisition

  • mods1r.20161207.0063-0066

  • mods2r.20161207.0071-0074

10:40 we send the science binocular script for G191B2B. Seeing 2.3", the clouds are getting to the target. The seeing keeps oscillating

  • mods1r.20161207.0067-0069

  • mods1b.20161207.0042-0044

  • mods2r.20161207.0075-0077

  • mods2b.20161207.0051-0053

10:50 we want to slew to the Lick star of Pancino, but it is in the direction of the wind, we cannot observe it now

10:51 the sky is completely covered with clouds. The TO closes the dome

11:57 the TO opens the dome

12:06 we send the acquisition of MODS Annibali program DD068_runB. We use new MOS alignment tool mosAlign_exp

12:13 the TO is having trouble with the pointing

12:15 we send the acquisition again

12:17 the right side (MODS2) got stuck

12:18 we send the acquisition again: error on MODS1, we have to abort the script

12:22 seeing is 1.4"

12:23 we send the acquisition again. Seeing is 0.9"

  • mods1r.20161207.0070-0072

  • mods2r.20161207.0078-0080

12:39 we send the science script. Seeing is 0.9". Out of spec in ParAngle

  • mods1r.20161207.0073-0075

  • mods1b.20161207.0045-0047

  • mods2r.20161207.0081-0083

  • mods2b.20161207.0054-0056

12:49 seeing is 0.7"-0.8"

13:06 we are in twilight, we stop the integration after 2 exposures. The last 1 images are BAD

13:07 we send the acquisition of Annibali standard star Feige34. Seeing about 1".

13:11 the TO needs to check the pointing. We abort the script

13:13 we send the acquisition again

  • mods1r.20161207.0076-0080
  • mods2r.20161207.0081-0086

13:27 we send the science of Feige34

  • mods1r.20161207.0081

  • mods1b.20161207.0048

  • mods2r.20161207.0087

  • mods2b.20161207.0057

13:40 we put MODS1 and MODS2 to sleep

-- %USERSIG{AlidaMarchetti - 2016-12-07}%

Comments

 
Topic revision: r7 - 20 Feb 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback